# “What If My Face Gets Scanned Without Consent”: Understanding Older Adults’ Experiences with Biometric Payment

### v1

@@ Abstract.
Biometric payment, i.e., biometric authentication implemented in digital payment systems, can reduce memory demands and streamline payment for older adults. However, older adults’ perceptions and practices regarding biometric payment remain underexplored. We conducted semi-structured interviews with 22 Chinese older adults, including both users and non-users. Participants were motivated to use biometric payment due to convenience and perceived security. However, they also worried about loss of control due to its password-free nature and expressed concerns about biometric data security. Participants also identified desired features for biometric payment, such as lightweight and context-aware cognitive confirmation mechanisms to enhance user control. Based on these findings, we outline recommendations for more controllable and informative digital financial services that better support older adults.
biometric payment, older adults
copyright: acmcopyrightjournalyear: 2021doi: $X$.XXXX/XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Biometric payment has gained global traction. The global facial recognition payment (FRP) market was valued at \$4.62 billion in 2022 (Research, 2023). Apple Pay employs biometric authentication as its default payment method (Apple, 2024). India’s BHIM Aadhaar Pay enables transactions via fingerprint verification linked to a national ID (BANKIT, ). In China, biometric payment is widely adopted in both online transactions through smartphones (Jao, 2019) and in offline environments such as convenience stores, supermarkets, restaurants, and hotels via self-checkout kiosks (Zhong et al., 2021; Liu et al., 2021; Gao et al., 2025). Fig. 1 illustrates the biometric payment process at a supermarket checkout kiosk in China.
@@ 2.1. Biometric Payment Usage
Overall, biometric payment offers both benefits, such as enhanced convenience (Zhong and Moon, 2022) and security (Yu et al., 2024), but also triggers concerns, such as security and privacy risks (Lee and Pan, 2023; Zhang and Zhang, 2024). However, existing research has predominantly focused on the general users, often drawing samples from younger generations. This study contributes an empirical investigation into older adults’ perceptions, experiences, and practices regarding biometric payment, providing insights into biometric payment use among this largely understudied population.
@@ 2.2. Older Adults and Digital Payment
Overall, older adults face unique challenges and hold distinct perceptions regarding digital payments. As a form of digital payment, biometric payment may involve different perceptions and practices, which remain underexplored. To address this gap, this paper explores older adults’ motivations, concerns, protective behaviors, and desired improvements related to biometric payment systems.
@@ 2.3. Older Adults and Biometric Technology
Several studies have also explored older adults’ perceptions and usage of biometric technologies (Kumalasari et al., 2024; Ahmed et al., 2017; Fenuku et al., 2024). For instance, Kumalasari et al. examined biometric self-authentication for pension fund withdrawals by older adults and found that facilitating conditions and behavioral intention to use were key drivers of continued use, while health-related limitations in vision, hearing, and mobility negatively affected their intention to use (Kumalasari et al., 2024). Ahmed et al. investigated the use of biometrics as a substitute for passwords and reported that most older adults perceived biometric options as more secure and providing a better experience than traditional password-based systems (Ahmed et al., 2017).
@@ 3.1. Participant Recruitment
Recruitment channels. Initially, we intended to hand out flyers, but we discovered that many older adults were hesitant to discuss sensitive topics, particularly those related to payment. Therefore, instead of directly distributing flyers, we visited places where older adults frequently gather such as senior activity centers and senior residential communities—a common approach to engaging with seniors (Hornung et al., 2017)—and started casual conversations to introduce ourselves, explain the purpose of our research, and build rapport. Additionally, to broaden our recruitment scope, we advertised our study through posters on social media platforms widely used by the older adult population in China, such as Douyin and WeChat groups. Prior to advertising, we obtained permission from group moderators. In our advertisement poster, we included the contact information of the first author. Since filling out a screening survey might have been time-consuming and burdensome for older adults, we assessed their eligibility during casual conversations, whether conducted in person or online. During the casual conversations, we primarily focused on confirming the participant’s age and whether they had used mobile payment, along with a few background questions such as which payment methods they use. This screening took only a few minutes. The online and offline procedures were the same, with online screening conducted via calls or text and offline screening conducted in person.
@@
Demographics. Table 1 describes the background of the interviewed participants. Their ages ranged from 50 to 73. Seventeen participants are women and five are men, perhaps because women were more willing to share their views on biometric payment. They came from seven cities spanning four tiers (i.e., Tier 1: Guangzhou and Shenzhen, New Tier 1: Chongqing and Chengdu, Tier 3: Handan, Tier 4: Jilin and Beihai). Among all 22 participants, 13 had used biometric payment, and the remaining had not. The main biometric payment methods utilized by older adults were face recognition (nn==$10$) and fingerprint recognition (nn==99). The majority of biometric payment terminals used were smartphones (nn==$13$). Among them, many mentioned employing facial or fingerprint recognition to complete online shopping purchases and bank transfers. Three participants also recounted using facial recognition at checkout terminals in physical stores. One older adult mentioned using facial recognition for automatic fare deduction when entering the subway.
@@ 3.2. Semi-structured Interview
Interview procedure. At the beginning of the interview, we clarified the research goal, the interview process and duration, potential benefits and risks, and how the interview data would be used. Participants could ask any questions about the study, and verbal informed consent was then obtained. Although we had an interview guide, the interviews were conducted in a semi-structured format, allowing flexibility to follow up on specific questions or topics in depth to elicit unexpected but relevant responses. Additionally, we skipped some questions if the participant had already addressed the information in their earlier answers. At the end of the interview, we encouraged the participant to provide any additional insights we might have overlooked.
@@ 3.3. Coding and Analysis
Since our coding process involved multiple iterative rounds of coding, discussion, and refinement to generate themes, codes here served as a process rather than a final product. This is one of the reasons for not seeking inter-rater reliability (IRR) according to McDonald et al. (McDonald et al., 2019). As such, we did not calculate IRR. All quotes were initially translated from Chinese to English using Google Translate. The first author then refined the translations to ensure accuracy.
@@ 3.4. Ethical Considerations
We followed our university’s Policy on the Ethical Conduct of Research Involving Human Participants, which is grounded in three overarching ethical principles: beneficence, justice, and respect for persons. Our research received approval from our institution’s Institutional Review Board (IRB). All audio recordings were transcribed, and the resulting transcripts were securely stored without any personal identifiers. Each participant’s transcript was labeled with generic identifiers such as “P01” and access to the data was strictly limited to the research team. Participation was entirely voluntary. Prior to obtaining consent, participants were thoroughly informed about the study’s purpose, data use, potential benefits and risks, and could ask any questions they were concerned about. Participants were informed that they could decline to participate without any repercussions. They were further assured of their right to skip any question for any reason or withdraw from the interview at any point without impacting the overall process. To help alleviate stress or concerns about difficulty in answering, we reassured participants that the research questions were focused on their personal thoughts and everyday experiences and that there were no right or wrong answers.
@@ 4. Findings
In this section, we present seniors’ motivations for using biometric payment in Sec. 4.1, including security, convenience, social influence, and facilitating conditions. Following that, Sec. 4.2 captures four types of concerns, including lack of control, security and privacy concerns, lack of knowledge about how to use biometric payment, and difficulties adapting to different payment scenarios. We also examine older adults’ protective measures when using biometric payment in Sec. 4.3. Lastly, we present our findings regarding older adults’ desired improvements for biometric payment in Sec. 4.4.
@@ 4.1. Motivation for Using Biometric Payment
Furthermore, even when participants are aware of the potential risks of biometric payment, they are willing to accept them in exchange for convenience. P15 (F, 65, user) reflected on this trade-off,
@@
One key motivation for some older adults to use biometric payment is the perception of enhanced security. They believe that biometric methods effectively prevent password exposure during transactions—a significant advantage over traditional password-based systems. P20 (M, 51, user) pointed out,
@@
Beyond recommendations from children, seeing people nearby use biometric payments also influences older adults to adopt the technology. As P18 (F, 50, user) shared, “I didn’t really overthink it. Many people were using it, so it just felt natural to set up…I saw how convenient it was for others, figured it was safe, and decided to go for it.”
@@ 5. Discussion
Based on the findings, we discuss three tensions in older adults’ experiences with biometric payment (Sec. 5.1), compare our findings with prior work on older adults’ usage of general digital payment and biometric technology (Sec. 5.2), and provide design implications for more inclusive biometric payment systems for older adults (Sec. 5.3).
@@ 5.1. Unraveling Biometric Payment Usage through the Lens of Older Adults
Tension between ease of use and security and privacy concerns. In Sec. 4.1, we found that even when some older adults were aware of the potential security and privacy risks associated with biometric payment, they still chose to use these systems for their perceived convenience. Prior research has identified a risk-benefit trade-off regarding facial recognition payment in the general population (Hu et al., 2023; Liu et al., 2021). Our study extends this finding to older adults, suggesting that their decision-making also reflects a similar calculus.
@@
Moreover, this tension aligns with findings from the general Chinese consumer market. For instance, Liu et al. found that in the early stages of facial recognition payment, many Chinese users were willing to accept privacy risks in exchange for perceived benefits (Liu et al., 2021). Cultural variations play a significant role in shaping this risk-benefit assessment. For example, this perspective contrasts with findings from studies conducted in the U.S., where biometric privacy is often taken more seriously. Kugler noted that many American users are so concerned about biometric privacy that they are even willing to forgo tangible benefits to protect their privacy (Kugler, 2019). One possible explanation is that Chinese users tend to prioritize convenience and functionality over privacy compared to users in Western countries, as shown in prior studies on technology use (Lowry et al., 2011; Liu et al., 2021; Cao and Everard, 2008), and their collectivist culture makes them more cooperative when responding to new technological trends (Johnston et al., 2009). Future research can explore potential cultural differences in this risk-benefit tension and how it might affect subsequent protective behaviors.
@@
Moreover, it remains uncertain whether the concerns expressed by older adults in Sec. 4.2 accurately reflect the actual risks involved or if some of these concerns are potentially overestimated. For instance, due to the experience with A4-sized printed photos potentially spoofing facial recognition machines at the workplace, one older adult assumed that the same level of security risk applied to biometric payment systems, as shown in Sec. 4.2.3. This perception might not align with the actual technological architecture and risk profiles of these systems, suggesting a potential gap between user perception and reality. Additionally, we observed in Sec. 4.3 that some older adults avoid video calls to prevent facial data leakage. However, without expert validation, it is difficult to determine whether these strategies are genuinely effective, necessary, or merely based on misconceptions about the risks. Therefore, future work should integrate expert perspectives to assess the legitimacy of older adults’ concerns and the effectiveness of their protective behaviors.
@@ 5.2. Positioning Biometric Payment in Older Adults’ Broader Technology Use
Nevertheless, biometric payment introduces additional perceived vulnerabilities and amplifies existing digital payment concerns. The perceived uncertainty of the technology and intangibility of provided services have been identified as issues in older adults’ interactions with mobile payment (Yang et al., 2015). The use of biometric authentication without requiring passwords may further intensify older adults’ perceived lack of control over the transaction process, as shown in Sec. 4.2.2. Similarly, while concerns around security and privacy have long existed in digital payment (Omotayo and Akinyode, 2020; Jin and Fan, 2022), biometric payment introduces additional and biometric-based concerns, such as biometric data leakage and hacking via biometric spoofing in Sec. 4.2.3. Future research aiming to understand biometric payment or guide system design should recognize these unique and biometric-specific aspects to better support older adults and promote more inclusive digital financial experiences. It is also worth noting that older adults may have a preference for one biometric payment method over another (e.g., some favored fingerprint payment over facial recognition payment because its deliberate action gives them a sense of control in Sec. 4.2.2). However, we did not conduct a systematic comparison due to the limited number of participants who have used multiple biometric payment methods. We call for future work to systematically examine user perceptions across different biometric payment methods.
@@ 5.3. Toward Inclusive Biometric Payment
Moreover, older adults’ expected cognitive confirmation mechanisms have several key characteristics. First, they tend to feature low cognitive load, suggesting that the design should strike a balance between security and usability and ensure that the cognitive burden remains manageable. Second, many of these mechanisms incorporate personalization. However, although personalized gestures and visual patterns strengthen users’ sense of control, this approach introduces new usability challenges, as personalized codes might need regular updates and reliable recall. The implementation needs to understand how often users are willing to change these inputs, how often such patterns must be updated, whether users can manage this effort, and how to effectively integrate biometric recognition with personalized codes to achieve a balance between low cognitive load and a strong sense of control. A third characteristic is context awareness. It can reduce the risk of unintended transactions by linking the verification process directly to the specific context of each transaction such as the transaction amount and items. However, it may raise privacy concerns in public settings about the leakage of transaction information. In addition, these methods also need to account for the diverse physical and cognitive abilities of older adults. For example, gesture-based contextual inputs might be more accessible for some users, while voice-based confirmations might be preferable for those with limited mobility or dexterity.
@@
Supporting older adults’ autonomy in adoption. Older adults have voiced a desire for both operational and informational guidance to enhance their autonomy in using biometric payment systems in Sec. 4.4. Prior research suggests that older adults tend to be more sensitive to loss aversion (Jin and Fan, 2022; Kahneman, 2011) and often face greater difficulties when using new technologies compared to younger users (Jiang et al., 2016; Choudrie et al., 2018; Yusif et al., 2016). As a result, providing sufficient operational guidance is crucial for building their usage confidence and empowering their autonomy.

### v2

# “What If My Face Gets Scanned Without Consent”: Older Adults’ Experiences with Biometric Payment

@@ Abstract.
Biometric payment, i.e., biometric authentication implemented in digital payment systems, can reduce memory demands and streamline payment for older adults. However, older adults’ perceptions and practices regarding biometric payment remain underexplored. We conducted semi-structured interviews with 22 Chinese older adults, including both users and non-users. Participants were motivated to use biometric payment due to convenience and perceived security. However, they also worried about loss of control due to its password-free nature and expressed concerns about biometric data security. Participants also identified desired features for biometric payment, such as lightweight and context-aware cognitive confirmation mechanisms to enhance user control. We outline recommendations for more accessible and informative digital financial services that better support older adults.
biometric payment, older adults, facial recognition payment, fingerprint recognition payment
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791562isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Biometric payment has gained global traction. The global facial recognition payment (FRP) market was valued at \$4.62 billion in 2022 (Research, 2023). Apple Pay employs biometric authentication as its default payment method (Apple, 2024). India’s BHIM Aadhaar Pay enables transactions via fingerprint verification linked to a national ID (BANKIT, ). In China, biometric payment is widely adopted in both online transactions through smartphones (Jao, 2019) and in offline environments such as convenience stores, supermarkets, restaurants, and hotels via self-checkout kiosks (Zhong et al., 2021; Liu et al., 2021; Gao et al., 2025). Figure 1 illustrates the biometric payment process at a supermarket checkout kiosk in China.
@@
Figure 1: The figure illustrates the biometric payment process at a supermarket self-checkout kiosk, showing four steps: payment initiation, facial recognition, payment confirmation, and payment success.
@@ 2.1. Biometric Payment Usage
Overall, biometric payment offers both benefits, such as enhanced convenience (Zhong and Moon, 2022) and security (Yu et al., 2024), but also triggers concerns, such as security and privacy risks (Lee and Pan, 2023; Zhang and Zhang, 2024). However, existing research has predominantly drawn insights from younger generations. Our study contributes an empirical investigation into older adults’ perceptions, experiences, and practices regarding biometric payment, providing insights into biometric payment use among this largely understudied population.
@@ 2.2. Older Adults and Digital Payment
Overall, older adults face unique challenges and hold distinct perceptions regarding digital payments. To what extent they carry over to biometric digital payment remains underexplored. To address this gap, this paper explores older adults’ motivations, concerns, protective behaviors, and desired improvements related to biometric payment systems.
@@ 2.3. Older Adults and Biometric Technology
Several studies have also explored older adults’ perceptions and usage of biometric technologies (Kumalasari et al., 2024; Ahmed et al., 2017; Fenuku et al., 2024). For instance, Kumalasari et al. examined biometric self-authentication for pension fund withdrawals by older adults and found that facilitating conditions and behavioral intention to use were key drivers of continued use, while health-related limitations in vision, hearing, and mobility negatively affected their intention to use (Kumalasari et al., 2024). Ahmed et al. investigated biometrics as a substitute for passwords and reported that most older adults perceived biometric options as more secure and providing a better experience than traditional password systems (Ahmed et al., 2017).
@@ 3.1. Participant Recruitment
Recruitment channels. Initially, we intended to hand out flyers, but we discovered that many older adults were hesitant to discuss sensitive topics, particularly those related to payment. Therefore, instead of directly distributing flyers, we visited places where older adults frequently gather such as senior activity centers and senior residential communities, a commonly used approach for engaging seniors (Hornung et al., 2017). We started casual conversations with them to introduce ourselves, explain the purpose and value of the study, and outline participation expectations, thereby inviting individuals to take part in the research and establishing rapport. Additionally, to broaden our recruitment scope, we advertised our study through posters on social media platforms widely used by the older adult population in China, such as Douyin and WeChat groups. Prior to advertising, we obtained permission from group moderators. In our advertisement poster, we included the contact information of the first author. Since filling out a screening survey might have been time-consuming and burdensome for older adults, we assessed their eligibility during casual conversations, whether conducted in person or online. During the casual conversations, we primarily focused on confirming the participant’s age and whether they had used mobile payment, along with a few background questions such as which payment methods they use. This screening took only a few minutes. The online and offline procedures were the same, with online screening conducted via calls or text and offline screening conducted in person.
@@
Demographics. Table 1 describes the background of the interviewed participants. Their ages ranged from 50 to 73. Seventeen participants are women and five are men. They came from seven cities spanning four tiers (i.e., Tier 1: Guangzhou and Shenzhen, New Tier 1: Chongqing and Chengdu, Tier 3: Handan, Tier 4: Jilin and Beihai). Among all 22 participants, 13 had used biometric payment, and the remaining had not. The main biometric payment methods utilized by older adults were face recognition (nn==$10$) and fingerprint recognition (nn==99). The majority of biometric payment terminals used were smartphones (nn==$13$). Among them, many mentioned employing facial or fingerprint recognition to complete online shopping purchases and bank transfers. Three participants also recounted using facial recognition at checkout terminals in physical stores. One older adult mentioned using facial recognition for automatic fare deduction when entering the subway.
@@
Table 1: The table presents demographic information of interviewed seniors, including ID, age, sex, city, bio-payment methods (face or fingerprint), and devices used (such as smartphone, checkout terminal, or fare gate).
@@ 3.2. Semi-structured Interview
Figure 2: The figure shows a flowchart of the interview process, starting with introduction and consent, followed by sections on basic understanding, perceptions, motivations, concerns, protective behaviors, desired improvements, and ending with feedback. Paths are shown for both users and non-users of biometric payment.
@@
Interview procedure. At the beginning of the interview, we showed the consent form with participants and further explained it such as the research goal, the interview process and duration, potential benefits and risks, and how the interview data would be used. Participants could ask any questions about the study, and audio-recorded oral consent was then obtained. [^2] According to guidance from our university’s Institutional Review Board (IRB), when conducting research where seeking written consent is not practical or too sensitive, audio-recorded oral consent might be less of a privacy risk than written consent and can be considered as an alternative. In our study, older adult participants are particularly sensitive to signing formal documents, especially in discussions related to payment. Although we had an interview guide, the interviews were conducted in a semi-structured format, allowing flexibility to follow up on specific questions or topics in depth to elicit unexpected but relevant responses. Additionally, we skipped some questions if the participant had already addressed the information in their earlier answers. At the end of the interview, we encouraged the participant to provide any additional insights we might have overlooked.
@@ 3.3. Coding and Analysis
Since our coding process involved multiple iterative rounds of coding, discussion, and refinement to generate themes, codes here served as a process rather than a final product, which is one of the reasons for not seeking inter-rater reliability (IRR) according to McDonald et al. (McDonald et al., 2019). All quotes were initially translated from Chinese to English using Google Translate. The first author then refined the translations to ensure accuracy.
@@ 3.4. Ethical Considerations
We followed our university’s Policy on the Ethical Conduct of Research Involving Human Participants, which is grounded in three overarching ethical principles: beneficence, justice, and respect for persons. [^4] https://vprd.hkust.edu.hk/policies-compliance/policies-guidelines/human-participants Our research received approval from our institution’s Institutional Review Board (IRB). All audio recordings were transcribed, and the resulting transcripts were securely stored without any personal identifiers. Each participant’s transcript was labeled with generic identifiers such as “P01” and access to the data was strictly limited to the research team. Participation was entirely voluntary. Prior to obtaining consent, participants were thoroughly informed about the study’s purpose, data use, potential benefits and risks, and could ask any questions they were concerned about. Participants were informed that they could decline to participate without any repercussions. They were further assured of their right to skip any question for any reason or withdraw from the interview at any point without impacting the overall process. To help alleviate stress or concerns about difficulty in answering, we reassured participants that the research questions were focused on their personal thoughts and everyday experiences and that there were no right or wrong answers.
@@ 4. Findings
In this section, we present seniors’ motivations for using biometric payment in Sec. 4.1, including security, convenience, social influence, and facilitating conditions. Following that, Sec. 4.2 captures four types of concerns, including lack of control, security and privacy concerns, lack of knowledge about how to use biometric payment, and difficulties adapting to different payment scenarios. We also examine older adults’ protective measures when using biometric payment in Sec. 4.3. Lastly, we present findings regarding their desired improvements for biometric payment in Sec. 4.4.
@@ 4.1. Motivation for Using Biometric Payment
Furthermore, even when participants are aware of the potential risks of biometric payment, they are still willing to accept them in exchange for convenience. P15 (F, 65, user) reflected on this trade-off,
@@
One key motivation is the perception of enhanced security. They believe that biometric methods effectively prevent password exposure during transactions—a significant advantage over traditional password-based systems. P20 (M, 51, user) mentioned,
@@
Beyond recommendations from children, seeing people nearby use biometric payments also influences older adults to adopt the technology. As P18 (F, 50, user) pointed out, “I didn’t really overthink it. Many people were using it, so it just felt natural to set up. I saw how convenient it was for others, figured it was safe, and decided to go for it.”
@@ 5. Discussion
Based on the findings, we discuss three tensions in older adults’ experiences with biometric payment (Sec. 5.1), compare our findings with prior work on older adults’ usage of general digital payment and biometric technology (Sec. 5.2), and provide design implications for more inclusive biometric payment systems (Sec. 5.3).
@@ 5.1. Unraveling Biometric Payment Usage through the Lens of Older Adults
Tension between ease of use and security and privacy concerns. In Sec. 4.1, we found that even when older adults were aware of the potential security and privacy risks associated with biometric payment, they still chose to use these systems for their perceived convenience. Prior research has identified a risk-benefit trade-off regarding facial recognition payment in the general population (Hu et al., 2023; Liu et al., 2021). Our study extends this finding to older adults, suggesting that their decision-making also reflects a similar calculus.
@@
Moreover, this tension aligns with findings from the general Chinese consumer market. For instance, Liu et al. found that in the early stages of facial recognition payment, many Chinese users were willing to accept privacy risks in exchange for perceived benefits (Liu et al., 2021). Cultural variations play a significant role in shaping this risk-benefit assessment. This perspective contrasts with findings from studies conducted in the U.S., where biometric privacy is often taken more seriously. Kugler noted that many American users are so concerned about biometric privacy that they are even willing to forgo tangible benefits to protect their privacy (Kugler, 2019). One possible explanation is that Chinese users tend to prioritize convenience and functionality over privacy compared to users in Western countries, as shown in prior studies on technology use (Lowry et al., 2011; Liu et al., 2021; Cao and Everard, 2008), and their collectivist culture makes them more cooperative when responding to new technological trends (Johnston et al., 2009). Future research can explore potential cultural differences in this risk-benefit tension and how it might affect subsequent protective behaviors.
@@
Moreover, it remains uncertain whether the concerns expressed by older adults in Sec. 4.2 accurately reflect the actual risks involved or if some of these concerns are potentially overestimated. For instance, due to the experience with A4-sized printed photos potentially spoofing facial recognition machines at the workplace, one older adult assumed that the same level of security risk applied to biometric payment systems (Sec. 4.2.3). This perception might not align with the actual technological architecture and risk profiles of these systems, suggesting a potential gap between user perception and reality. Additionally, we observed in Sec. 4.3 that some older adults avoid video calls to prevent facial data leakage. However, without expert validation, it is difficult to determine whether these strategies are genuinely effective, necessary, or merely based on misconceptions about the risks. Therefore, future work should integrate expert perspectives to assess the legitimacy of older adults’ concerns and the effectiveness of their protective behaviors.
@@ 5.2. Positioning Biometric Payment in Older Adults’ Broader Technology Use
Nevertheless, biometric payment introduces additional perceived vulnerabilities and amplifies existing digital payment concerns. The perceived uncertainty of the technology and intangibility of provided services have been identified as issues in older adults’ interactions with mobile payment (Yang et al., 2015). The use of biometric authentication without requiring passwords may further intensify older adults’ perceived lack of control over the transaction process, as shown in Sec. 4.2.2. Similarly, while concerns around security and privacy have long existed in digital payment (Omotayo and Akinyode, 2020; Jin and Fan, 2022), biometric payment introduces additional and biometric-based concerns, such as biometric data leakage and hacking via biometric spoofing in Sec. 4.2.3. Future research aiming to understand biometric payment or guide system design should recognize these unique and biometric-specific aspects to better support older adults and promote more inclusive digital financial experiences. It is also worth noting that older adults may have a preference for one biometric payment method over another (e.g., some favored fingerprint payment over facial recognition payment because its deliberate action gives them a sense of control, as shown in Sec. 4.2.2). However, we did not conduct a systematic comparison due to the limited number of participants who have used multiple biometric payment methods. We call for future work to systematically examine user perceptions across different biometric payment methods.
@@ 5.3. Toward Inclusive Biometric Payment
Moreover, older adults’ expected cognitive confirmation mechanisms have several key characteristics. First, they tend to feature low cognitive load, suggesting that the design should strike a balance between security and usability and ensure that the cognitive burden remains manageable. Second, many of these mechanisms incorporate personalization. While personalized gestures and visual patterns strengthen users’ sense of control, this approach introduces new usability challenges, as personalized codes might need regular updates and reliable recall. The implementation needs to understand how often users are willing to change these inputs, how often such patterns must be updated, whether users can manage this effort, and how to effectively integrate biometric recognition with personalized codes to achieve a balance between low cognitive load and a strong sense of control. A third characteristic is context awareness. It can reduce the risk of unintended transactions by linking the verification process directly to the specific context of each transaction such as the transaction amount and items. However, it may raise privacy concerns in public settings about the leakage of transaction information. In addition, these methods also need to account for the diverse physical and cognitive abilities of older adults. For example, gesture-based contextual inputs might be more accessible for some users, while voice-based confirmations might be preferable for those with limited mobility or dexterity.
@@
Supporting older adults’ autonomy in adoption. Older adults have voiced a desire for both operational and informational guidance to enhance their autonomy in using biometric payment systems (Sec. 4.4). Prior research suggests that older adults tend to be more sensitive to loss aversion (Jin and Fan, 2022; Kahneman, 2011) and often face greater difficulties when using new technologies compared to younger users (Jiang et al., 2016; Choudrie et al., 2018; Yusif et al., 2016). As such, providing sufficient operational guidance is crucial for building their usage confidence and autonomy.